Jill is 22.5 kilometers away from Joey. Both begin to walk towards each other at the same time. Jill walks at 1.5 km/hr. They meet in five hours. How fast is Joey walking?
a) 1.5 k/hr b) 2.25 km/hr c) 3 km/hr d) 6 km/hr e) 15 km/hr

Answer by richwmiller(17219)

d=22.5 km
Jill=1.5*5=7.5
Joey=x*5=15
x=3 km/h
check
1.5*5+3*5=22.5